3.)If we know the values of the sine and cosine of a and b, we can find the value of sin(A+B) by using the ____formula for sine. State the formula :




Sine(A+B)




4.)if we know the values of the sine and cosine A ana B, we can find the value of cos (A-B) by using the ____ formula for cosine. State the formula :




Cos(A-B)




5.)if we know the values of sin x and cos x, we can find the value of sin 2x by using the ____formula for sine. State the formula :sin 2x =____




6.)if we know the value of cos x and the quadrant in which x/2 lies, we can find the value of sin(x/2) by using the _____ formula for sine. State the formula:




Sin(x/2)=______

1
Expert's answer
2022-01-20T11:49:56-0500
  1. Formula sum of angles: "sin(A + B) = sin A cos B + cos A sin B"
  2. Formula difference of angles: "cos(\u03b1\u2212\u03b2)=cos(\u03b1)\u22c5cos(\u03b2)+sin(\u03b1)\u22c5sin(\u03b2)"
  3. double angle formula: "sin(2x)=2sin(x)\u22c5cos(x)"
  4. half-angle formula: "\sin(\frac{x}{2})=±\sqrt{1-cos(x)\over 2}"

+ (positive) if "{x \over 2}" lies in quadrant | or ||

- (negative) if "{x \over 2}"   lies in quadrant ||| or |V